# Complexity of Einstein-Maxwell-non-minimal coupling $R^{2}F^{2}$: the role of the penalty factor

### v1

@@ Abstract
In this paper, we examine the holographic complexity of a theory of Einstein-Maxwell with a non-minimal coupling $R^{2}F_{\mu\nu}F^{\mu\nu}$ term via complexity=anything. We introduce a perturbative black brane solution in AdS spacetime up to the first order of the non-minimal coupling. Because the model exhibits the resistivity proportional to the temperature, it could describe strange metals. The complexity of the solution could be characterized by three quantities: the conserved charge, the non-minimal coupling and the generalized term in the complexity. The generalized term is chosen to be $C^{2}$, the Weyl squared tensor, $R^{2}F^{2}$ and $F^{2}$. The physical interpretation of the three quantities has been provided where the generalized term is interpreted as a bulk penalty factor, the non-minimal coupling and the conserved charge as an effective scrambling time of the dual theory which contains the choice of the generalization parameter.
@@ 1 Introduction
The AdS/CFT (Anti-de Sitter/Conformal Field Theory) duality provides a powerful framework for understanding the relationship between gravitational theories in asymptotically AdS spaces and their corresponding quantum field theories on the boundary [1]. Although, our universe is a dS space, studying the duality in AdS is valuable to find a deep insights to develop the duality to dS/CFT and real-world physics [2, 3, 4, 5, 6, 7, 8]. This duality suggests that complex gravitational phenomena can be analyzed through the lens of quantum information theory, enabling the exploration of measures such as the quantum computational complexity. The complexity can be viewed as a measure of how difficult it is to prepare a given quantum state from a reference state using a sequence of operations [9]. The complexity serves as a crucial indicator of the degrees of freedom [10, 11] and the scrambling dynamics in a quantum system, revealing insights into how information is processed and stored in black hole backgrounds. By examining complexity within the AdS/CFT framework, we gain valuable insights into the structure of quantum states and their holographic properties, ultimately bridging the gap between deep gravitational insights and operational aspects of quantum field theories.
@@
Recently, it has been proposed that there is a generalization in the complexity and the volume of an arbitrary region of spacetime [15].
@@
Non-minimal coupling in gravity, particularly when invoking the AdS (Anti-de Sitter) framework, has gained significant attention due to its potential to enhance our understanding of gravitational dynamics in various physical contexts. Due to the curvature coupling, these non-minimal theories have a significant effect on the higher curvature regimes such as near black hole or the early universe. The non-minimal term could be viewed as deforming the boundary theory, which means that n-point functions and transport coefficients would be altered. Specifically, this term breaks the translational symmetry which is more appropriate to model realistic crystals where containing impurity [32]. In addition, the non-minimal theories could model strange metals where the resistivity scales linearly with temperature [33]. The model we have considered in this paper represents a linear relation between the resistivity and the temperature which signaling the behavior of strange metals [34].
The organization of the paper is as follows: we introduce a perturbative solution for a non-minimal coupling $R^{2}F^{2}$ of Einstein-Maxwell theory in the section 2. Subsequently, the complexity growth rate of the solution in the complexity equals anything proposal is considered in the section 3. The complexity growth rate in this solution is dependent to the conserved charge $Q$, the non-minimal coupling $q_{2}$ and the term of the generalization $\gamma b$. There are some choices for the generalization of the complexity where we have chosen the Weyl squared tensor $\gamma C^{2}$, $\gamma F^{2}$ and $\gamma R^{2}F^{2}$. Then, in the section 4 we provide the physical interpretations of the conserved charge as a kind of restriction on the simple unitary operators, the non-minimal coupling as the effective scrambling time of the quantum circuit and the choice of the generalization in the quantum circuit as the penalty factor in the quantum circuit. At the end of the section 4 we provide two examples in superconducting circuits that qualitatively fit with the penalty factor interpretation and finally we conclude in the section 5.
@@ 2 Non-minimal $R^{2}F_{\mu\nu}F^{\mu\nu}$ AdS Black Brane Solution
at large curvature regime, when $4q_{2}R^{2}\gg q_{1}$ the coefficient of the kinetic term in the Maxwell equation flips sign which means the ghost instability. Then, the sign of the coupling constant $q_{2}$ should be negative. Considering that the spacetime in our model is 4-dimensional with a two-dimensional spatial direction, we adopt the following ansatz for the metric:
@@
We will use this ansatz to solve Eq. (10), where the potential 1-form is expressed as:
@@
The solution of $h(r)$ is as:
@@
Since the equations for the tttt–component and the $rr$-component are identical up to first order, we conclude that $8f_{0}(r)r^{3}H_{0}^{\prime}(r)=0$. Therefore, we determine that $H_{0}$ is given as follows:
@@
We impose the condition that Eq. (A) and Eq. (A) must be equivalent up to the first order of $q_{2}$. Consequently, $H_{1}$ is calculated as follows:
@@ 3 Holographic Complexity of $R^{2}F^{2}$
The growth rate of the generalized complexity can be considered by the behavior of the conserved momentum in full-time. The late-time behavior of the growth rate is given by:
@@
The numerical analysis of the $\mathcal{C}_{Any}$-boundary time diagrams for the Weyl squared tensor $a^{2}=C^{2}$ are plotted in Fig.(1) and the complexity-boundary time in Fig.(2), for $R^{2}F^{2}$ are Fig.s (7), (4) and for $F^{2}$ in Fig.s(5), (6). Regarding the fact that the complexity being defined as a maximum value, the intersections of the complexity-boundary time diagrams indicate points where complexity switches from one upper branch to another, known as $\tau_{turning}$. At this time, a transition occurs in $\mathcal{C}_{Any}$-boundary time diagrams.
Figure 1: CGR in terms of the boundary time for $b=C^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
@@
Figure 3: CGR in terms of the boundary time for $b=R^{2}F^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
@@
Figure 5: CGR in terms of the boundary time for $b=F^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
@@
4 Physical interpretation of CGR
As Fig.s(2), (4) and (6) show, by adopting different parameters of $Q$, $q_{2}$ and $\gamma$, the diagrams differ in the locations of the branches and the distance between the branches. The main question is that if the differences in the diagrams of CGR have a physical origin or not. In [38], it shows that the locations of the late time CGR (where the branches are divided) are controlled by the value of the bulk fields in the case $b=C^{2}$. For example, if there is an scalar field in the bulk, at the locations of the late time CGR, $r_{i}$, there is the following relation [38]:
$\Big(\nabla_{i}<T_{jk}>\Big)^{2}\sim\partial_{i}\phi^{(0)}\partial^{i}\phi^{(0)}=\frac{1-g^{(0)}_{tt}}{2\gamma g^{(0)}_{tt}}$, (51)
where $T_{jk}$ is the energy-momentum tensor at $r_{i}$, the location of the late time of CGR, $\phi^{(0)}$ is the value of the scalar field at the cut-off $r_{i}$ and $g_{tt}^{(0)}$ is the $tt$-component of the metric (12) in Fefferman-Graham coordinates at the cut-off $r_{i}$.
The same reasoning holds for the other choices of $b$.
@@ 4.1 The generalization term $\gamma b$
In this section the generalization term is interpreted as the penalty factor in the boundary circuit.
Let us rewrite (40) as follows:
@@
The above relation is similar to Nielson geometrical method to compute complexity in the quantum theory [39]:
where the metric is defined in a Hilbert space of the unitary operators:
$ds^{2}=\sum_{\alpha}\gamma_{\alpha}(Y^{\alpha})^{2}dt^{2}.$ (54)
To make the cost function, $G_{IJ}$ in (53) acting differently in directions, there introduces a penalty factor which weighting some gates heavily and some lightly by a coefficient into the cost function. In the way of mere similarity, it seems that the coefficient $(r^{d-1}a)^{2}$ in (52) would be a position-dependent penalty factor or let us call it curvature penalty factor.
The mere similarity could be broaden by the bulk symplectic form and Berry curvature form in gauge/gravity duality. Let us starting with $\mathcal{C}_{QFT}$ and the duality of the variation of gravitational observables (which is the bulk symplectic form) and the Berry curvature form [40]. The Berry curvature form could be written in terms of the imaginary part of the quantum geometric tensor or Fubini-Study (FS) metric [41] in the boundary theory :
@@
$\Omega$ is the bulk symplectic form and $\omega$ the Berry curvature form. Furthermore, FS metric is related to $G_{IJ}$ as:
@@
where $\mathcal{P}_{IJ}$ is the penalty factor matrix. One could rewrite (53) in terms of the bulk symplectic form as follows:
@@
In the last line the factor $\mathcal{P}_{IM}\mathcal{P}_{NJ}\delta\int d\sigma\sqrt{g^{\prime}_{ij}\dot{x}^{i}\dot{x}^{j}}$ could be seen as a source for the penalty factor $\mathcal{P}(a)$, where $a$ is the generalization factor in (52). By change of $a$ the penalty factor in $\mathcal{C}_{QFT}$ is deformed.
As a result, by the change of $\gamma b$, $\mathcal{C}_{Any}$ and $\dot{\mathcal{C}}_{Any}$ are changed. A question may arise: is there a penalty factor in quantum circuits that its change makes the counterpart of $\mathcal{C}_{Any}$ as the gate complexity and $\dot{\mathcal{C}}_{Any}$ as the depth of the circuit [9] changed? we introduce two of these circuits in the section 4.4 which are applicable in quantum computations.

### v2

@@ Abstract
We investigate holographic complexity in Einstein-Maxwell theory with a non-minimal coupling of the form $R^{2}F_{\mu\nu}F^{\mu\nu}$ within the complexity=anything framework. A perturbative AdS black brane solution is constructed to first order in the non-minimal coupling parameter. Owing to the linear temperature dependence of the resistivity, this model provides a holographic realization of strange metal behavior. The complexity growth rate (CGR) is governed by three independent parameters: the conserved charge, the non-minimal coupling, and the choice of the generalized term entering the complexity functional. We consider three representative generalizations, namely the Weyl tensor squared, $R^{2}F^{2}$ , and $F^{2}$. We provide a physical interpretation of these parameters, the generalized bulk functional analytically induces a deformation of the effective cost metric, which can be interpreted as a bulk penalty factor, while the conserved charge and the non-minimal coupling control an effective scrambling time in the dual theory. The role of the generalization parameter is shown to be closely tied to the structure of the corresponding quantum circuit.
@@ 1 Introduction
The AdS/CFT (Anti-de Sitter/Conformal Field Theory) duality provides a powerful framework for understanding the relationship between gravitational theories in asymptotically AdS spaces and their corresponding quantum field theories on the boundary [1]. Although, our universe is a dS space, studying the duality in AdS is valuable to find a deep insights to develop the duality to dS/CFT and real-world physics [2, 3, 4, 5, 6, 7, 8, 9]. This duality suggests that complex gravitational phenomena can be analyzed through the lens of quantum information theory, enabling the exploration of measures such as the quantum computational complexity. The complexity can be viewed as a measure of how difficult it is to prepare a given quantum state from a reference state using a sequence of operations [10]. The complexity serves as a crucial indicator of the degrees of freedom [11, 12] and the scrambling dynamics in a quantum system, revealing insights into how information is processed and stored in black hole backgrounds. By examining complexity within the AdS/CFT framework, we gain valuable insights into the structure of quantum states and their holographic properties, ultimately bridging the gap between deep gravitational insights and operational aspects of quantum field theories.
@@
More recently, these ideas have been unified and extended through the complexity=anything conjecture [16], which allows complexity to be defined via a broad class of diffeomorphism-invariant bulk functionals. This generalization introduces an intrinsic ambiguity in the definition of complexity, encoded in a choice of scalar functional that weights different regions of spacetime.
@@
Non-minimal coupling in gravity, particularly when invoking the AdS (Anti-de Sitter) framework, has gained significant attention due to its potential to enhance our understanding of gravitational dynamics in various physical contexts. Due to the curvature coupling, these non-minimal theories have a significant effect on the higher curvature regimes such as near black hole or the early universe. The non-minimal term could be viewed as deforming the boundary theory, which means that n-point functions and transport coefficients would be altered. Specifically, this term breaks the translational symmetry which is more appropriate to model realistic crystals where containing impurity [33]. In addition, the non-minimal theories could model strange metals where the resistivity scales linearly with temperature [34]. The model we have considered in this paper represents a linear relation between the resistivity and the temperature, a hallmark of strange metals [35]. Motivated by this observation, we study a four-dimensional Einstein-Maxwell theory with a non-minimal coupling of the form $R^{2}F_{\mu\nu}F^{\mu\nu}$. Since exact solutions are not available, we construct a perturbative AdS black brane solution to first order in the non-minimal coupling.
The main goal of this work is to explore how such non-minimal interactions and the choices of the generalized parameter affect holographic complexity. Within the complexity=anything framework, we analyze the time dependence of the complexity growth rate (CGR) and show that it is controlled by three parameters: the conserved charge, the non-minimal coupling, and the choice of generalization in the complexity functional. We consider three representative choices, the Weyl tensor squared, $R^{2}F^{2}$, $F^{2}$, and provide a physical interpretation of their effects.
An important aspect of quantum circuit complexity is the freedom in choosing the set of elementary gates and the associated cost or penalty factors that weight different directions in the space of unitary transformations. In Nielsen’s geometric formulation, this freedom is encoded in the choice of a cost metric, and different penalty schemes correspond to inequivalent notions of complexity. In holographic settings, however, the microscopic identification of boundary gates remains intrinsically ambiguous [16]. As a result, holographic complexity proposals should be understood as defining equivalence classes of circuit complexities rather than unique microscopic constructions. This intrinsic ambiguity motivates the study of generalized complexity functionals and their geometric implications.
The three generalization choices allow us to systematically explore how different generalizations modify the effective geometry governing extremal bulk trajectories associated with complexity. We show analytically that the generalized functional deforms the effective cost metric, in a manner structurally equivalent to the introduction of penalty factors in Nielsen’s circuit complexity. This identification does not rely on specifying a unique set of boundary gates; rather, it emerges at the level of complexity geometry and reflects the inherent freedom in defining cost structures in holographic complexity.
We further examine the implications of this deformation for the time dependence of complexity growth. In particular, we analyze the CGR and identify an effective scrambling time that controls this behavior. We emphasize that this timescale should be understood as a complexity-dependent effective scrambling time, rather than the microscopic scrambling time defined via out-of-time-order correlators or shock-wave analyses. While the intrinsic chaotic properties of the bulk geometry remain unchanged, the effective scrambling time governing complexity growth is shown analytically to depend on both the non-minimal coupling and the generalized parameter through their influence on the effective complexity geometry.
The paper is organized as follows: in the section 2 we construct the perturbative black brane solution. In the section 3, we introduce the generalized complexity functional and analyze the resulting complexity growth for the three choices of the generalized parameter. The section 4 is devoted to the interpretation of the generalized functional as encoding a penalty structure and to the analysis of the effective scrambling time governing complexity growth and supporting this interpretations by presenting qualitative analogies with superconducting quantum circuits. We conclude in the section 5 with a discussion of our results and possible future directions.
@@ 2 Non-minimal $R^{2}F_{\mu\nu}F^{\mu\nu}$ AdS Black Brane Solution
at large curvature regime, when $4q_{2}R^{2}\gg q_{1}$ the coefficient of the kinetic term in the Maxwell equation flips sign which means the ghost instability (i.e., the spectrum is not bounded from below). Then, the sign of the coupling constant $q_{2}$ should be negative. Considering a four dimensional spacetime with a two-dimensional spatial subspace, we adopt the following ansatz for the metric:
@@
We will use this ansatz to solve Eq. (10), where the potential 1-form is expressed as  [^1] We have chosen the radial gauge $A_{r}=0$ and assumed translational symmetry in the boundary directions, which together with staticity leads to the vanishing of the spatial components. Furthermore, to ensure the regularity of the gauge potential at the future horizon, we have imposed the condition that $h(r)$ vanishes at the horizon $r=r_{h}$.:
@@
The solution of $h(r)$ is as  [^2] The solution is given by the solving (10) by use of ansatz (12) and (13).:
@@
Since the equations for the tttt–component and the $rr$-component are identical up to first order, we conclude that $8f_{0}(r)r^{3}H_{0}^{\prime}(r)=0$. Therefore, we determine that $H_{0}$ is given as follows  [^3] By the condition $8f_{0}(r)r^{3}H^{\prime}_{0}(r)=0$ we deduce that $H_{0}(r)$ is a constant. We apply this condition that the speed of light on the boundary is equal to one. Therefore, $H_{0}(r)=0$.:
@@
By subtraction of A from A and solving it, $H_{1}$ is calculated as follows:
@@ 3 Holographic Complexity of $R^{2}F^{2}$
The time evolution of the generalized complexity growth rate is governed by the behavior of the conserved momentum $P_{v}$. The late-time behavior of the growth rate is given by:
@@
We now explain that the numerical analysis shows the relationship between the boundary time $\tau$ and the conserved momentum $P_{v}$ (which sets the complexity growth rate, $\frac{d\mathcal{C}_{Any}}{d\tau}\propto P_{v}$) for the cases $b$ equals to $C^{2}$, $R^{2}F^{2}$ and $F^{2}$. The panels demonstrate how this relationship changes with different parameters ($Q$, $q_{2}$ and $\gamma$). Regarding the fact that the complexity being defined as a maximum value, the intersections of the complexity-boundary time diagrams indicate points where complexity switches from one upper branch to another, known as $\tau_{turning}$  [^4] The different branches in the CGR arises from $\frac{d\mathcal{C}_{Any}}{d\tau}\sim P_{v}\sim\sqrt{\tilde{U}}$. The CGR is a multivalued function of $r_{min}$, then there are two branches of the two extremal surfaces (with the same boundary condition and anchoring), but the complexity computed from one of them at some critical time dominates as the same transition is shown in [39].. At this time, a transition occurs in $\mathcal{C}_{Any}$-boundary time diagrams.
Figure 1: The CGR in terms of the boundary time for $b=C^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
@@
Figure 3: The CGR in terms of the boundary time for $b=R^{2}F^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
@@
Figure 5: The CGR in terms of the boundary time for $b=F^{2}$ (a) with the parameters Q=0.09Q=-0.09, $q_{2}=-0.06$ and $\gamma=-0.002$. (b) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=-0.002$. (c) Q=0.09Q=-0.09, $q_{2}=-0.03$ and $\gamma=0.002$. (d) Q=0.1Q=-0.1, $q_{2}=-0.03$ and $\gamma=-0.002$.
@@
4 Physical interpretation of The CGR
As shown in Figs.(2), (4) and (6), varying the parameters $Q$, $q_{2}$ and $\gamma$ leads to noticeable differences in the structure of the complexity growth rate (CGR). In particular, both the locations of the branches and the separation between them depend sensitively on these parameters. A natural question is whether these differences admit a clear physical interpretation.
It was shown in [40] that, for the choice $b=C^{2}$, the late time values of the CGR, namely the points at which different branches merge, are controlled by the dynamics of the bulk energy-momentum tensor. For instance, one finds at the location $r_{i}$ corresponding to the late time of the CGR:
$2\gamma g^{(0)}_{tt}\mathscr{C}^{2}=r_{i}^{2}-g^{(4)}_{tt}$. (51)
where $\mathscr{C}$ is the Weyl squared tensor evaluated at $r_{i}$, the location of the late time of the CGR, $g^{(4)}_{tt}\sim<T_{tt}>$, for $T_{tt}$ the component of the energy-momentum tensor and $g_{tt}^{(0)}$ is the $tt$-component of the metric (12) in Fefferman-Graham coordinates at the cut-off $r_{i}$. An analogous reasoning applies to the other choices of the generalization parameter $b$.
In the following subsections, we provide a physical interpretation of the dependence of the CGR on each of the three parameters $\gamma b$, $q_{2}$, and $Q$.
@@ 4.1 The generalization term $\gamma b$
We begin by interpreting the generalized term $\gamma b$ as a penalty factor in the dual quantum circuit.
In holographic approaches to quantum complexity, the complexity of a boundary state is computed by extremizing a bulk functional, which defines a geometric structure governing the optimal trajectories associated with state preparation. In Nielsen’s geometric formulation of circuit complexity, this structure is encoded in a cost metric on the space of unitary operators, where penalty factors weight different directions corresponding to distinct classes of quantum gates [41]:
where the metric is defined in a Hilbert space of unitary operators:
$ds^{2}=\sum_{\alpha}\gamma_{\alpha}(Y^{\alpha})^{2}dt^{2}.$ (53)
Here, the coefficients $\gamma_{\alpha}$ act as penalty factors, weighting different directions in the space of gates and suppressing costly operations.
Within the complexity=anything framework, the generalized scalar functional entering the bulk action (40) modifies the effective geometry in which the extremal surface or particle trajectory is evaluated:
@@
where $g^{\prime}_{ij}=r^{2(d-1)}a^{2}g_{ij}$. As shown in (54), the generalized term enters multiplicatively in the integrand and therefore deforms the metric governing the bulk trajectory. This deformation is fully analogous to the introduction of penalty factors in circuit complexity, where certain directions in the space of unitaries are assigned a higher computational cost.
Importantly, this identification does not rely on specifying a unique set of boundary gates or a microscopic circuit construction, which is known to be intrinsically ambiguous in holographic duality [16]. Rather, the generalized functional encodes a penalty structure at the level of the effective complexity geometry. Different choices of the generalized parameter correspond to inequivalent cost metrics and therefore to different notions of optimal circuit complexity, all of which are consistent within the inherent freedom of defining holographic complexity.
This interpretation is further supported by rewriting the complexity functional in a form directly comparable to Nielsen’s action, where the generalized factor appears as a position-dependent weight in the effective metric, the factor $(r^{d-1}a)^{2}$ in (54), which we refer to as a curvature penalty factor. In this sense, the generalized term provides a bulk realization of penalty factors that control the relative cost of different bulk trajectories, and consequently modify the resulting CGR.
We emphasize that this penalty-factor interpretation should be understood as a structural correspondence rather than a microscopic identification. The intrinsic ambiguity in the choice of gates and penalties in quantum circuits is mirrored holographically by the freedom to choose the generalized functional, and the analytical relation derived here makes this correspondence explicit.
This analogy can be strengthened using the correspondence between bulk symplectic forms and boundary Berry curvature. In holography, variations of bulk observables are related to the Berry curvature form $\omega$ in the dual theory [42], which is itself tied to the imaginary part of the quantum geometric tensor or Fubini–Study (FS) metric [43]:
@@
$\Omega$ is the bulk symplectic form and $\omega$ the Berry curvature form. The FS metric can be related to $G_{IJ}$ as:
@@
where $\mathcal{P}_{IJ}$ encodes the penalty factors. One could rewrite (52) in terms of the bulk symplectic form as follows:
@@
Expressing the circuit complexity in terms of the bulk functional then shows explicitly that variations of $a(r)$ source deformations of the penalty matrix $\mathcal{P}_{IJ}$. Consequently, modifying $\gamma b$ directly corresponds to deforming the penalty structure of the dual quantum circuit. Given that holography does not uniquely specify a gate set or penalty scheme, any gravitational complexity proposal necessarily corresponds to an equivalence class of quantum circuits. Within this framework, we show analytically that the generalized functional deforms the effective cost metric in a manner identical to the introduction of a penalty factor in Nielsen’s geometric approach. In this precise sense, the generalized term encodes a bulk realization of circuit penalties.
A natural question arises: is there a penalty factor in quantum circuits whose variation alters the corresponding quantities, namely $\mathcal{C}_{Any}$ as the gate complexity and $\dot{\mathcal{C}}_{Any}$ as the circuit depth [10]? In the section 4.4, we introduce two such circuit constructions that are applicable to quantum computation.